I have no clue how to solve 3t/4-t/2=1 can you help me?
Answer by checkley77(12844)

3t/4-t/2=1 the common denominator is 4.
(3t-2t)/4=1 cross multiply
3t-2t=4*1
t=4 ans.
Proof:
3*4/4-4/2=1
12/4-2=1
3-2=1
1=1